### v1

# Thermodynamic Gravity with Non-Extensive Horizon Entropy and Topological Calibration

Abstract
We revisit Jacobson’s thermodynamic derivation of gravitational dynamics in the presence of generalized, non-extensive horizon entropies. Working within a local Rindler-wedge framework, we formulate the Clausius relation as the stationarity condition of a Massieu functional at fixed Unruh temperature, which identifies the entropy slope as the parameter controlling the effective gravitational coupling. For area-type entropies with constant slope, the construction reproduces Einstein’s equations with $G_{\rm eff}=1/(4s_{0})$, while curvature-dependent entropy densities supplemented by an internal entropy-production term yield the field equations of $f(R)$ gravity.
Motivated by group-entropic considerations and long-range correlations, we model the entropy of horizon cross sections by a power law $S(A)=\eta(A/4G)^{\delta}$ and analyze its local and global implications. To fix the otherwise arbitrary coarse-graining scale entering the entropy slope, we introduce a Topological Calibration Principle that ties the reference area to intrinsic geometric data through the Gauss–Bonnet theorem. For compact two-dimensional sections, this selects a canonical calibration area and leads to a topology-dependent effective coupling $G_{\rm eff}(\chi)\propto|\chi|^{1-\delta}$ with $\chi$ representing the ensuing Euler characteristic. Consistency across scales and topologies yields logarithmic bounds on $|1-\delta|$, while the associated scale dependence induces a characteristic modulation of the gravitational coupling in cosmology. The framework thus provides a controlled route to confront non-extensive horizon thermodynamics with both theoretical consistency requirements and observational constraints.
Keywords:
emergent gravity, alternative gravity theories, horizon thermodynamics, non—extensive entropy
1 Introduction
Over the past two decades, the thermodynamic interpretation of gravity has evolved into a coherent organizing principle for spacetime dynamics. Stationary black holes behave as genuine thermodynamic systems, characterized by a temperature set by the surface gravity and an entropy proportional to the area of the event horizon Bekenstein (1973); Hawking (1975, 1976); Bardeen et al. (1973); Smarr (1973b, a). Within this framework, the Bekenstein–Hawking relation
$S_{\rm BH}\ =\ \frac{A}{4G}\,$, (1)
(with the horizon area $A$ and Newton’s gravitational constant $G$) is not merely a dimensional estimate; rather, it provides a bridge between classical general relativity, quantum field theory in curved spacetime, and statistical mechanics.
Jacobson’s seminal work sharpened this connection by showing that the Einstein field equations can be understood as an equation of state. Specifically, if each local Rindler horizon is assigned an entropy density proportional to its area together with an Unruh temperature, then imposing the Clausius relation $\delta Q=TdS$ for all local causal horizons yields the Einstein equations with a cosmological constant Jacobson (1995); Unruh (1976); Parikh and Svesko (2018). In this picture, the Einstein equations appear as macroscopic consistency conditions for a local horizon thermodynamic identity.
Such an emergent viewpoint is conceptually complementary to programs that treat the metric (or connection or vielbein) as fundamental variables subject to direct quantization. Instead of postulating gravitons as microscopic degrees of freedom, one seeks in this picture a thermodynamic underpinning for the spacetime metric itself, while remaining agnostic about the precise nature of the underlying constituents — whether they originate from vacuum entanglement across causal horizons Bombelli et al. (1986); Iorio et al. (2004); Jacobson (2016); Van Raamsdonk (2010), collective excitations of a pre-geometric medium, or more exotic structures. Closely related developments, notably those of Padmanabhan et al., highlight the role of gravitational heat densities, equipartition of degrees of freedom on holographic screens, and the extremization of appropriately defined entropy functionals Padmanabhan (2010, 2005). In these frameworks, spacetime dynamics follows from a variational principle applied to an entropy functional built from null vectors, with the Einstein field equations arising as consistency conditions for its extremum. By contrast, Verlinde’s entropic gravity proposal casts this logic in a manifestly holographic and information-theoretic language, interpreting Newtonian gravity — and, in suitable limits, the Einstein equations — as entropic forces associated with changes in information on screens endowed with area-scaling entropy Verlinde (2011). Despite differences in technical realization and in the interpretation of what is meant by “emergence,” these approaches share a key conceptual idea, namely that thermodynamic quantities such as entropy, temperature, and heat flux on horizons or screens are taken as primary, while the spacetime metric and its dynamics encode an effective macroscopic response.
On the other hand, semiclassical and many-body considerations (further fueled by results from AdS/CFT and AdS/CMT correspondence Zaanen et al. (2015)) indicate that the simple area law (1) is only the leading term of a more general entropic form. Entanglement entropy across a horizon generically exhibits ultraviolet divergences proportional to the area of the entangling surface, but also sub-leading corrections that depend on field content and background curvature for field-theoretic and holographic discussions Srednicki (1993); Solodukhin (2011); Rangamani and Takayanagi (2017); Casini et al. (2011); Bueno and Myers (2015); Bueno et al. (2015); Blanco et al. (2013). Wald’s Noether-charge construction Wald (1984, 1993) identifies, in a general diffeomorphism-invariant theory, an entropy functional which reduces to (1) in Einstein gravity but acquires higher-curvature and topological contributions in more general models Wald (1993); Iyer and Wald (1994). At the same time, systems with long-range interactions and strong correlations, including self-gravitating matter, often display non-additive thermodynamics and anomalous scaling of entropy with system size Padmanabhan (1990); Chavanis (2002); Campa et al. (2009). Motivated by these facts, several authors have recently proposed various non-extensive (or non-additive) generalizations of horizon entropy $S(A)$ Tsallis (1988); Tsallis and Cirto (2013); Caruso and Tsallis (2008); Jizba and Lambiase (2022); Moradpour et al. (2020); Saridakis et al. (2018); Abbasi and Gharaati (2020).
In gravitational applications, such generalized entropies are often implemented in a phenomenological way: one replaces $S_{\rm BH}$ by a modified entropy $S(A)$, inserts it into a first-law relation and reads off effective corrections to black-hole mechanics, to Friedmann equations or to entropic-force constructions. While such models are useful as parametrizations of deviations from general relativity, they leave open several conceptual questions from the emergent-gravity perspective. If gravitational field equations are to arise from horizon thermodynamics, a generalized entropy must be compatible with a local Clausius relation on Rindler patches and with a sensible identification of temperature, energy and effective gravitational coupling. In particular, the entropy slope $dS/dA$ should play a central role, as it controls both the macroscopic normalization of horizon entropy and the strength of the coupling in Jacobson-type constructions.
A recent framework proposed by Lü, Di Gennaro and Ong  et al. (2025) takes an important step in this direction. In particular, working in a local-horizon framework in the spirit of Jacobson, they replace the Bekenstein–Hawking entropy by a Tsallis–Jensen functional Tsallis and Jensen (2025) defined through a deformed Legendre structure, chosen so as to preserve a sensible canonical temperature on the Rindler patch. In the spherically symmetric sector the thermodynamic manifold is effectively one-dimensional, and the Clausius relation can be integrated to identify entropy and energy functionals under suitable regularity assumptions. The resulting equations of state describe an effective Newton coupling that becomes a function of area, together with emergent field equations that depart from the Einsteinian form while retaining a clear thermodynamic interpretation. Their analysis clarifies to what extent non-extensive entropies can be incorporated in an emergent-gravity setup without internal inconsistencies, but also makes manifest that many non-extensive models used in gravitational physics are, in effect, one-dimensional truncations where integrability constraints are relatively mild.
In this work, we follow a complementary and more conservative route. Rather than undertaking a full global analysis of horizon thermodynamics on a multi-parameter state space, we ask what can be inferred locally from a Jacobson-type construction once a non-extensive entropy slope is given. Concretely, we model the entropy of a horizon cross-section by the phenomenological power law $S(A)=\eta(A/4G)^{\delta}$, which is also known as Tsallis-$\delta$ entropy Tsallis and Cirto (2013), Barrow entropy Barrow (2020) or Barrow–Tsallis entropy Jizba et al. (2024). In its microcanonical form it also coincides with Tsallis–Jensen entropy Tsallis and Jensen (2025) which belong to the class of the so-called group entropies Tempesta (2016b, 2011, a); Jensen and Tempesta (2024). In the following, we will simply use the name power law entropy as it will better fit our purpose. With the power law entropy at hand, we will analyze how the corresponding entropy slope feeds into the effective Newton constant on local Rindler wedges.
To conceptually underpin our construction, we introduce a Topological Calibration Principle (TCP), which links the reference area used to evaluate the slope to the intrinsic geometry of horizon cross-sections via the Gauss-Bonnet theorem. In two dimensions, the Gauss-Bonnet theorem provides the unique relation among area, intrinsic curvature, and Euler characteristic, so employing it to set the intrinsic calibration scale allows to implement a “no external scales” requirement within an emergent-gravity setting, and promotes topology to a quantitative criterion constraining admissible deviations from extensivity.
In summary, we use horizon thermodynamics primarily as a consistency framework for generalized entropy laws, rather than as a model-building device. On the local side, a canonical Jacobson-like construction identifies the entropy slope as the unique macroscopic parameter that fixes the effective coupling in the field equations, reproducing Einstein gravity in the area-type branch and providing a non-equilibrium route to $f(R)$ theories when $s(x)\propto f^{\prime}(R)$. On the global side, the TCP uses intrinsic curvature and Euler characteristic to calibrate the resolution area, turning $(\eta,\delta)$ into parameters that control both the normalization and the scale and topology dependence of $G_{\rm eff}$. Taken together, these ingredients lead to explicit logarithmic bounds on $|1-\delta|$ and to structural constraints on non-extensive horizon entropies. These bounds become physically meaningful once one assumes a universal pair $(\eta,\delta)$ and identifies the effective coupling inferred from horizon thermodynamics with the low-energy Newton constant. Within this set of assumptions, topology becomes a falsifiable diagnostic for viable departures from the Bekenstein–Hawking area law in emergent-gravity scenarios.
The paper is organized as follows. In Sec. 2 we recast Jacobson’s local horizon thermodynamics in canonical form and show that, for area-type entropies, the resulting field equations reduce to Einstein’s equations with $G_{\rm eff}=1/(4s_{0})$, while curvature-dependent entropy densities reproduce $f(R)$ gravity in a non-equilibrium setting. Sec. 3 introduces the Topological Calibration Principle, which fixes the evaluation scale $A_{*}$ through the Gauss–Bonnet relation and translates the generalized entropy slopes into topology- and scale-dependent couplings $G_{\rm eff}(\chi,A_{*})$, leading to explicit bounds on $|1-\delta|$. Finally, in Sec. 4, we summarize the main results and discuss the implications for both microscopic models and phenomenology. For the reader’s convenience, technical details related to the Gauss–Bonnet entropy are collected in the Appendix.
2 Field equations from local horizon thermodynamics
A central motivation for exploring generalized horizon entropies is the tension between holographic area laws and standard thermodynamic extensivity. For three-dimensional systems with short–range interactions, ordinary Boltzmann–Gibbs statistical mechanics yields an extensive entropy $S_{\rm BG}\propto V\sim L^{3}$ whenever the number of configurations grows exponentially with the number of constituents, $W(N)\sim\exp(cN)$. Black–hole horizons and many quantum systems with strong correlations instead display an area law, $S\propto A\sim L^{2}$, which is non–extensive in this sense Amico et al. (2008); Caruso and Tsallis (2008). In strongly gravitating systems, this mismatch is accompanied by familiar pathologies of self–gravitating ensembles — inequivalence of microcanonical and canonical descriptions, negative heat capacities and long–lived quasi–stationary states — signalling that the Boltzmann–Gibbs framework can become problematic Padmanabhan (1990); Chavanis (2002); Campa et al. (2009).
In this context, non–additive entropies have been advocated as effective macroscopic descriptions for long-range interacting and gravitational systems, including black–hole thermodynamics and cosmology Caruso and Tsallis (2008); Tsallis and Cirto (2013); Saridakis et al. (2018). A particularly economical class is provided by Tsallis–Jensen group entropies, which preserve the first three Shannon–Khinchin axioms while incorporating a nontrivial composability structure in Tempesta’s sense Tempesta (2011, 2016a); Jensen and Tempesta (2024); Tsallis and Jensen (2025). Concretely, one considers functionals of the form
$S_{\alpha,\gamma}[p]\ =\ \lambda\,\big[S_{R}^{(\alpha)}[p]\big]^{1/\gamma}\,$, (2)
where $\lambda$ is a constant, $\gamma>0$ is a deformation index, and
$S_{R}^{(\alpha)}[p]\ \equiv\ \frac{1}{1-\alpha}\,\ln\!\Big(\sum_{i}p_{i}^{\alpha}\Big)\,$, (3)
is the Rényi entropy. For a microcanonical ensemble with $W(A)$ equiprobable configurations, labeled by $i=1,\dots,W(A)$, one has $p_{i}=1/W(A)$, so that
$S_{R}^{(\alpha)}[p_{\rm eq}]\$$=$ $\ \frac{1}{1-\alpha}\,\ln\!\Bigg(\sum_{i=1}^{W(A)}p_{i}^{\alpha}\Bigg)\ =\ \frac{1}{1-\alpha}\,\ln\!\Big[W(A)\big(W(A)^{-1}\big)^{\alpha}\Big]$ (4)
$=$ $\ \frac{1}{1-\alpha}\,\ln\big(W(A)^{1-\alpha}\big)\ =\ \ln W(A)\,$,
and therefore
$S_{\alpha,\gamma}(A)\ \propto\ [\ln W(A)]^{1/\gamma}\,.$ (5)
In this sense, a group-entropic functional adapted to long-range, strongly correlated systems, when evaluated on a microcanonical horizon ensemble, quite generically produces power-law deformations of the Bekenstein–Hawking area law. To make this more specific, suppose that the number of microstates associated with a horizon cross-section grows in a stretched-exponential way Hanel and Thurner (2011),
[…]
so that
$\ln W(A)\ \propto\ \Big(\frac{A}{4G}\Big)^{q}\,.$ (7)
Evaluating the Tsallis–Jensen functional (2) on this microcanonical ensemble then yields
$S_{\alpha,\gamma}(A)\ \propto\ \Big[\ln W(A)\Big]^{1/\gamma}\ \propto\ \Big(\frac{A}{4G}\Big)^{q/\gamma}\,.$ (8)
In this framework the ratio $q/\gamma$ plays the role of an effective non-extensivity index, while $\lambda$ and numerical factors can be absorbed into an overall normalization.
Guided by these considerations, and without committing to a specific underlying microscopic model, we adopt the phenomenological power-law constitutive relation
$S(A)\ =\ \eta\,\Big(\frac{A}{4G}\Big)^{\delta},\qquad\eta\ >\ 0,\quad\delta\ >\ 0\,$, (9)
with $\eta$ and $\delta$ playing the role of effective entropic indices that encode possible departures from the Bekenstein–Hawking law. In the units $c=\hbar=k_{B}=1$ used throughout, the combination $A/4G$ is dimensionless, so that $S$ is dimensionless and both $\eta$ and $\delta$ are pure numbers. This makes (9) a dimensionally consistent deformation of the Bekenstein–Hawking relation (1). The Bekenstein–Hawking case corresponds to $\eta=1$ and $\delta=1$, while $\delta\neq 1$ describes non–extensive deformations of the area law in the sense discussed above. In a schematic microcanonical picture one may formally identify $\delta=q/\gamma$ and absorb proportionality constants into $\eta$, but in what follows we treat $(\eta,\delta)$ as phenomenological parameters to be constrained by consistency conditions rather than fixed by a specific microstate counting.
For concreteness, we will often illustrate the discussion in the window $\tfrac{1}{2}<\delta\leq 1$, corresponding to mild departures from the area law and to the regime that will be most relevant for the phenomenological bounds below. The canonical construction itself only assumes that $S(A)$ is differentiable and monotonic, hence the formal steps apply to any $\delta>0$; correspondingly, our constraints will be expressed in terms of $|1-\delta|$ and therefore cover both branches $\delta<1$ and $\delta>1$. As a useful benchmark on the non–extensive side, group–entropic microcanonical counting for area-law (or stretched–exponential) phase–space growth in $d$ spatial dimensions typically yields an effective scaling $S(A)\propto A^{\delta}$ with $\delta=d/(d-1)$, giving $\delta=3/2$ for $d=3$ Caruso and Tsallis (2008); Tsallis and Cirto (2013); Tsallis and Jensen (2025). Such values provide a theoretically motivated reference point against which the emergent-gravity consistency conditions and the observed near constancy of $G_{\rm eff}$ can be confronted, the latter favouring exponents close to the Bekenstein–Hawking case $\delta=1$.
With this phenomenological horizon law in place, we now turn to its local consequences. The central question is how a given entropy slope
$s_{\delta}(A_{*})\ \equiv\ \frac{dS}{dA}\Big|_{A_{*}}\,$, (10)
which quantifies the response of horizon entropy to area variations at a resolution scale $A_{*}$, translates into an effective gravitational coupling on local Rindler patches. In particular, for the non-extensive entropy law
$S(A)\ =\ \eta\Big(\frac{A}{4G}\Big)^{\delta}\,,\qquad\eta\ >\ 0\,,\quad\tfrac{1}{2}\ <\ \delta\ \leq\ 1\,$, (11)
one has
$s_{\delta}(A_{*})\ =\ \eta\,\delta\,(4G)^{-\delta}\,A_{*}^{\delta-1}\,$, (12)
and the canonical construction below will reproduce
$G_{\rm eff}(A_{*})\ =\ \frac{1}{4\,s_{\delta}(A_{*})}\ =\ \frac{G}{\eta\,\delta}\left(\frac{4G}{A_{*}}\right)^{\delta-1}\,$, (13)
in agreement with the running-$G$ picture of Ref.  et al. (2025).
We work on a smooth, time-oriented Lorentzian manifold $(\mathcal{M},g_{ab})$ with Levi-Civita connection and curvature conventions
$R^{a}{}_{bcd}\ =\ \partial_{c}\Gamma^{a}_{bd}\ -\ \partial_{d}\Gamma^{a}_{bc}\ +\ \Gamma^{a}_{ce}\,\Gamma^{e}_{bd}\ -\ \Gamma^{a}_{de}\,\Gamma^{e}_{bc}\,$,
$R_{bd}\ =\ R^{a}{}_{bad}\,,\qquad R\ =\ g^{bd}R_{bd}\,.$ (14)
Metric variations are taken with respect to the contravariant components $g^{ab}$, and we use the shorthand
$\delta_{g}F\ \equiv\ \left.\frac{d}{d\epsilon}F[g(\epsilon)]\right|_{\epsilon=0}\,,\qquad\delta g^{ab}\ \equiv\ \left.\frac{d}{d\epsilon}g^{ab}(\epsilon)\right|_{\epsilon=0}\,$, (15)
for a one-parameter family $g_{ab}(\epsilon)$ with $g_{ab}(0)=g_{ab}$.
To implement the local construction, we first recall the geometry of a local Rindler wedge. Around any spacetime point $p$ one can construct a wedge whose boundary $\mathcal{H}$ is generated by an affinely parametrized null congruence $k^{a}$ with affine parameter $\lambda$ such that $\lambda=0$ at $\Sigma$. In Jacobson’s setup we consider the portion of $\mathcal{H}$ to the past of $\Sigma$, hence $\lambda\in[-\lambda_{c},0]$ with $\lambda_{c}>0$. We choose a spacelike two-surface $\Sigma$ through $p$ as a local bifurcation surface, and pick a wedge-preserving boost Killing field $\chi^{a}$ that vanishes on $\Sigma$ and is tangent and null on $\mathcal{H}$. With this normalization, near $\Sigma$ one has
$\chi^{a}\ =\ -\,\kappa\,\lambda\,k^{a}\ +\ \mathcal{O}(\lambda^{2}),\qquad T\ =\ \frac{\kappa}{2\pi}\,,\quad\beta\ =\ \frac{2\pi}{\kappa}\,$, (16)
where $T$ is the local Unruh temperature associated with the boost flow and $\beta$ the corresponding inverse temperature Unruh (1976). The hypersurface element on $\mathcal{H}$ is
$d\Sigma^{b}\ =\ k^{b}\ d\lambda\ dA\,$, (17)
with $dA$ the induced area element on the spacelike two-dimensional cross-sections of $\mathcal{H}$ at fixed affine parameter $\lambda$.
Figure 1: Local Rindler-wedge geometry near a spacetime point $p$ (with $p\in\Sigma$). The boundary $\mathcal{H}$ is a null hypersurface generated by an affinely parametrized null congruence $k^{a}$, emanating from the spacelike local bifurcation two-surface $\Sigma$. An auxiliary null vector $\ell^{a}$ fixes the transverse direction, and $dA$ denotes the induced area element on the spacelike two-dimensional cuts of $\mathcal{H}$.
To describe the focusing of the null generators, we introduce an auxiliary null vector $\ell^{a}$ with $k^{a}\ell_{a}=-1$ and the projector onto the screen space
$h_{ab}\ =\ g_{ab}\ +\ k_{a}\ell_{b}\ +\ \ell_{a}k_{b}\,.$ (18)
The optical tensor
$B_{ab}\ \equiv\ h_{a}{}^{c}h_{b}{}^{d}\nabla_{c}k_{d}\,$, (19)
decomposes into expansion, shear and twist as
$B_{ab}\ =\ \tfrac{1}{2}\theta\,h_{ab}\ +\ \sigma_{ab}\ +\ \omega_{ab}\,$, (20)
where $\theta\equiv B^{a}{}_{a}$, $\sigma_{ab}$ is the trace-free symmetric part and $\omega_{ab}$ the antisymmetric part. Since $k^{a}$ is chosen to generate the null hypersurface $\mathcal{H}$ (hence it is hypersurface-orthogonal), Frobenius’ theorem implies the integrability condition $k_{[a}\nabla_{b}k_{c]}=0$, and therefore the twist (vorticity) tensor vanishes, $\omega_{ab}=0$ (see e.g. Wald (1984); Poisson (2004)).
For an affine congruence $k^{b}\nabla_{b}k^{a}=0$ the Raychaudhuri equation reads
$\frac{d\theta}{d\lambda}\ =\ -\,\tfrac{1}{2}\ \!\theta^{2}\ -\ \sigma_{ab}\sigma^{ab}\ -\ R_{ab}\,k^{a}k^{b}\,$, (21)
while the shear obeys the Sachs equation, projected on the screen space
$\frac{d\sigma_{ab}}{d\lambda}\ =\ -\ \!\theta\,\sigma_{ab}\ -\ \mathcal{C}_{ab}\ -\ \Big(\sigma_{ac}\sigma^{c}{}_{b}\ -\ \tfrac{1}{2}h_{ab}\,\sigma_{cd}\sigma^{cd}\Big)\,$, (22)
where $\mathcal{C}_{ab}$ denotes the screen-projected (trace-free) tidal term along $k^{a}$, expressed in terms of the Weyl tensor $C_{abcd}$ as
$\mathcal{C}_{ab}\ \equiv\ h_{a}{}^{c}h_{b}{}^{d}C_{ecfd}\,k^{e}k^{f}\,.$ (23)
By construction, $\mathcal{C}_{ab}$ is symmetric and trace–free with respect to $h_{ab}$, so the right–hand side of (22) is itself trace–free and therefore compatible with the definition of $\sigma_{ab}$ as the trace–free part of the optical tensor $B_{ab}$.
Imposing the initial conditions Jacobson (1995)
$\theta|_{\Sigma}\ =\ 0\,,\qquad\sigma_{ab}|_{\Sigma}\ =\ 0\,$, (24)
and assuming that the congruence is affinely parametrized and that $R_{ab}$ and the optical data $(\theta,\sigma_{ab},\omega_{ab})$ are smooth in a neighbourhood of $p$, Eq. (22) implies $\sigma_{ab}=O(|\lambda|)$ near $\lambda=0$, so that $\sigma_{ab}\sigma^{ab}=\mathcal{O}(\lambda^{2})$. Feeding this into (21) shows that
$\theta^{2}\;\;\;\mbox{and}\;\;\;\sigma_{ab}\sigma^{ab}\ =\ \mathcal{O}(\lambda^{2})\,$, (25)
and hence, to first non-trivial order in $\lambda$,
$\frac{d\theta}{d\lambda}\ =\ -\,R_{ab}k^{a}k^{b}\ +\ \mathcal{O}(|\lambda|)\,.$ (26)
Expanding $\theta(\lambda)$ near $\Sigma$ for sufficiently small affine cutoff $\lambda_{c}$ (i.e. for $|\lambda|\leq\lambda_{c}$)
$\theta(\lambda)\ =\ \theta|_{\Sigma}\ +\ \lambda\Big.\frac{d\theta}{d\lambda}\Big|_{\Sigma}\ +\ \mathcal{O}(\lambda^{2})\,$, (27)
and using $\theta|_{\Sigma}=0$ gives
$\theta(\lambda)\ =\ -\,\lambda\,R_{ab}k^{a}k^{b}\ +\ \mathcal{O}(\lambda^{2})\,.$ (28)
Since the area element on the cuts changes along the generators according to the standard definition of the expansion
$\theta\ \equiv\ \frac{1}{dA}\,\frac{d(dA)}{d\lambda}\quad\Rightarrow\quad\frac{d(dA)}{d\lambda}\ =\ \theta\ dA\,$, (29)
Eq. (28) implies, to first non-trivial order in $\lambda$
$\frac{d(dA)}{d\lambda}\ =\ -\,\lambda\,R_{ab}k^{a}k^{b}\,dA\ +\ \mathcal{O}(\lambda^{2})\,.$ (30)
Integrating along the generators over the past segment $\lambda\in[-\lambda_{c},0]$ then yields the linear area response of the horizon patch
$\delta_{g}\!\Big(\!\int_{\mathcal{H}}dA\Big)\ =\ -\!\int_{-\lambda_{c}}^{0}\!\lambda\,R_{ab}k^{a}k^{b}\,d\lambda\,dA\ +\ \mathcal{O}(\lambda_{c}^{3})\,$, (31)
so the leading (linear-in-$\lambda$) variation of the area element along the generators is controlled entirely by the Ricci focusing term $R_{ab}k^{a}k^{b}$, and the corresponding integrated area change over $\lambda\in[-\lambda_{c},0]$ scales as $\mathcal{O}(\lambda_{c}^{2})$. The quadratic optical terms $\theta^{2}$ and $\sigma_{ab}\sigma^{ab}$ enter Raychaudhuri at $\mathcal{O}(\lambda^{2})$ and therefore correct $\theta(\lambda)$ only at $\mathcal{O}(\lambda^{3})$, yielding contributions to the integrated area change that are higher order (starting at $\mathcal{O}(\lambda_{c}^{4})$) and consistently negligible at the order retained.
In this setup, the boost field $\chi^{a}$ also plays the role of modular generator Casini et al. (2017). The Bisognano–Wichmann theorem implies that, in Minkowski spacetime, the vacuum state reduced to a Rindler wedge is thermal at inverse temperature $\beta=2\pi/\kappa$ with respect to the boost charge Bisognano and Wichmann (1975, 1976)
[…]
Using the near-horizon expansion (16) and $d\Sigma^{b}=k^{b}\,d\lambda\,dA$, one finds, to $\mathcal{O}(\lambda_{c}^{2})$
$K\ =\ -\,\kappa\!\int_{-\lambda_{c}}^{0}\!\lambda\,T_{ab}k^{a}k^{b}\,d\lambda\,dA\ +\ \mathcal{O}(\lambda_{c}^{3})\,$, (33)
so that the reduced matter state on the wedge is canonically thermal, $\rho\propto e^{-\beta K}$ with $\beta=2\pi/\kappa$.
To phrase the local dynamics in canonical form it is convenient to introduce the Massieu (also known as Legendre) functional
$\Psi[g;\beta]\ =\ S_{G}[g]\ -\ \beta\,\langle K\rangle_{\rho}\,$, (34)
where $S_{G}[g]$ is the gravitational (horizon) entropy functional and $\langle K\rangle_{\rho}$ the expectation value of $K$ in the fixed matter state $\rho$. In the present canonical setting we vary only the geometry at fixed inverse temperature and matter state
$\delta\beta\ =\ 0\,,\qquad\delta\rho\ =\ 0\,$, (35)
and require stationarity with respect to patch-localized metric variations:
$\delta_{g}\Psi\ =\ 0\qquad\Leftrightarrow\qquad\delta_{g}S_{G}\ =\ \beta\,\delta_{g}\langle K\rangle_{\rho}\,.$ (36)
This implements local canonical equilibrium under $\delta\beta=\delta\rho=0$. We choose the variation to preserve, to leading order, the affine parametrization of $k^{a}$, the normalization $k\!\,\cdot\,\!\ell=-1$ and the initial conditions $\theta|_{\Sigma}=\sigma_{ab}|_{\Sigma}=0$. In particular, we require that the boost Killing field in the varied geometry $g_{ab}(\epsilon)$ be normalized in the same way as in Eq. (16), namely
$\chi^{a}(\epsilon)\ =\ -\,\kappa\,\lambda\,k^{a}(\epsilon)\ +\ \mathcal{O}(\lambda^{2})\,$, (37)
with the same surface gravity $\kappa$. Differentiating at $\epsilon=0$ one gets
$\delta\chi^{a}\ \equiv\ \left.\frac{d\chi^{a}(\epsilon)}{d\epsilon}\right|_{\epsilon=0}\ =\ \mathcal{O}(\lambda^{2})\,$, (38)
so that the additional contribution of $\delta\chi^{a}$ to $\delta_{g}\langle K\rangle$ scales as $\int_{-\lambda_{c}}^{0}\lambda^{2}\,d\lambda=O(\lambda_{c}^{3})$ and is sub-leading with respect to the $O(\lambda_{c}^{2})$ terms retained in the Raychaudhuri expansion.
In what follows we use the shorthand
$\delta_{g}\langle K\rangle\ \equiv\ \left.\frac{d}{d\epsilon}\,\langle K[g(\epsilon)]\rangle\right|_{\epsilon=0}\,$, (39)
for variations of the boost energy with respect to the geometry at fixed matter state $\rho$.
In the area-type branch we assume that, to leading order on the patch, the entropy density is constant,
$s(x)\ \equiv\ s_{0}\,$, (40)
with $s_{0}$ identified with the local slope of a global entropy $S(A)$,
$s_{0}\ \equiv\ \frac{dS}{dA}\Big|_{A_{*}}\,$, (41)
where $A_{*}$ is the coarse-graining area scale. In this case
$S_{G}[g]\ =\ \int_{\mathcal{H}}s_{0}\,dA\,,\qquad\delta_{g}S_{G}\ =\ s_{0}\,\delta\!\!\int_{\mathcal{H}}\!dA\,$, (42)
and Eq. (31) gives
$\delta_{g}S_{G}\ =\ -\,s_{0}\!\int_{-\lambda_{c}}^{0}\!\lambda\,R_{ab}k^{a}k^{b}\,d\lambda\,dA\ +\ \mathcal{O}(\lambda_{c}^{3})\,.$ (43)
On the “heat” side, at fixed $\beta$, the variation of the boost charge provides the heat flux across the patch. Using (33) together with the expansion (37) and the same $\mathcal{O}(\lambda_{c})$ counting as above, we evaluate $\delta_{g}\langle K\rangle$ at fixed matter state $\rho$, so that possible metric–induced changes of $T_{ab}$ contribute only beyond the $\mathcal{O}(\lambda_{c}^{2})$ order retained [^1] Operationally, $\delta\rho=0$ means that the local excitation content of the wedge is held fixed while varying the geometry of the patch; any induced response of $T_{ab}$ affects $\beta\,\delta_{g}\langle K\rangle$ only at higher order in the small-$\lambda$ expansion, and therefore does not enter the leading null-projected balance.. One then finds
$\beta\,\delta_{g}\langle K\rangle\ =\ -\,2\pi\!\int_{-\lambda_{c}}^{0}\!\lambda\,T_{ab}k^{a}k^{b}\,d\lambda\,dA\ +\ \mathcal{O}(\lambda_{c}^{3})\,.$ (44)
With the convention $\lambda\in[-\lambda_{c},0]$ and $k^{a}$ future-directed, the integrand $\lambda\,T_{ab}k^{a}k^{b}$ is non-negative when $T_{ab}k^{a}k^{b}\geq 0$, so positive energy flux corresponds to positive heat into the wedge. Eq. (44) is the boost-energy flux across the patch at fixed $\beta$.
Putting together the entropy and heat variations, and inserting (43) and (44) into the stationarity condition (36), one finds, to $\mathcal{O}(\lambda_{c}^{2})$
$s_{0}\!\int_{-\lambda_{c}}^{0}\!\lambda\,R_{ab}k^{a}k^{b}\,d\lambda\,dA\ =\ -\,2\pi\!\int_{-\lambda_{c}}^{0}\!\lambda\,T_{ab}k^{a}k^{b}\,d\lambda\,dA\ +\ \mathcal{O}(\lambda_{c}^{3})\,.$ (45)
The common factor $\int_{-\lambda_{c}}^{0}\lambda\,d\lambda=-\lambda_{c}^{2}/2$ is nonzero (and the same on both sides) and therefore divides out on both sides, while the $O(\lambda_{c}^{3})$ corrections are consistently neglected at the order kept in the Raychaudhuri expansion. One is thus left with the pointwise null identity
$s_{0}\,R_{ab}k^{a}k^{b}\ =\ 2\pi\,T_{ab}k^{a}k^{b}\;\;\;\;\text{for every null }\;\;\;\;k^{a}\text{ at }p\,$, (46)
which is the key local statement: stationarity of the Massieu functional ties the null-projected Ricci focusing to the null energy flux, and it holds for every null direction $k^{a}$ at $p$. The overall sign follows from our choice of orientation on the horizon generators and from the convention that positive $T_{ab}k^{a}k^{b}$ corresponds to positive energy flow into the wedge. A simultaneous reversal $k^{a}\!\to\!-k^{a}$ with the corresponding affine reparametrization leaves (46) invariant, so the ensuing tensor equation is independent of this convention.
To make the ensuing tensor structure explicit, it is convenient to define
$H_{ab}\ \equiv\ s_{0}\,G_{ab}\ -\ 2\pi\,T_{ab}\,$, (47)
and to recall that, for null vectors, $G_{ab}k^{a}k^{b}=R_{ab}k^{a}k^{b}$. Equation (46) then implies $H_{ab}k^{a}k^{b}=0$ for all null $k^{a}$. A standard algebraic lemma in Lorentzian geometry (see e.g. Wald (1984)) states that any symmetric tensor with vanishing double contraction on all null vectors must be proportional to the metric [^2] A short proof is obtained by working in a local orthonormal frame and using linear combinations of null vectors to show that all eigenvalues of $H^{a}{}_{b}$ coincide, so that $H_{ab}=\Phi\,g_{ab}$ for some scalar $\Phi$.
$H_{ab}\ =\ \Phi\,g_{ab}\,$, (48)
for some scalar function $\Phi$. Taking the covariant divergence and using $\nabla^{a}G_{ab}=0=\nabla^{a}T_{ab}$ gives $\nabla_{b}\Phi=0$, hence $\Phi=\Lambda$ is a constant. Dividing by $s_{0}$ one finally arrives at the Einstein field equations
$G_{ab}\ =\ 8\pi\,G_{\rm eff}\,T_{ab}\ +\ \Lambda g_{ab}\,,\qquad G_{\rm eff}\ =\ \frac{1}{4\,s_{0}}\,.$ (49)
In the area–type branch, the canonical stationarity condition for $\Psi=S_{G}-\beta\langle K\rangle$ on local Rindler patches closes onto Einstein’s equations, with a Newton coupling fixed by the leading entropy response on the patch. When the entropy density is constant at this order, $S_{G}=\int_{\mathcal{H}}s_{0}\,dA$, the entire gravitational input of the thermodynamic balance reduces to the single number $s_{0}=dS/dA|_{A_{*}}$, and the effective coupling follows immediately as $G_{\rm eff}=1/(4s_{0})$.
It is useful at this point to clarify how restrictive this area-type assumption really is. In the derivation above $s(x)$ was taken to be constant along the generators at leading order, so that only its global value $s_{0}$ appeared in the Clausius balance. If instead one allows $s(x)$ to vary already at order $\mathcal{O}(\lambda^{0})$, the structure of the thermodynamic identity changes qualitatively and naturally accommodates modified gravity theories. A minimal and well-studied example is obtained by letting the entropy density depend on the scalar curvature as
$s(x)\ =\ \eta\,f^{\prime}\!\big(R(x)\big)\,,\;\;\;\;\eta>0\ \;\;\;\text{constant}\,$, (50)
a choice directly motivated by Wald’s Noether-charge formula for $f(R)$ theories, in which the horizon entropy density is proportional to $f^{\prime}(R)$ Wald (1993); Iyer and Wald (1994). In other words, $s(x)$ is taken to coincide with the local density entering Wald’s entropy for a purely $f(R)$ Lagrangian. [^3] See Wald (1993); Iyer and Wald (1994) for the derivation from the action functional; here we only use that, for $f(R)$ theories, the Wald entropy density associated with a horizon cross-section is proportional to $f^{\prime}(R)$. With this choice one has generically $k\!\cdot\!\nabla s=\eta\,k\!\cdot\!\nabla f^{\prime}(R)\neq 0$ already at $\lambda=0$, so the local Clausius relation
$\delta Q\ =\ T\,\delta S\,$, (51)
is no longer integrable on the patch: the entropy change associated with a given geometric perturbation depends on the microscopic path followed along the generators. Following the non-equilibrium analysis of Eling, Guedens and Jacobson Eling et al. (2006), one introduces an internal production term $d_{i}S$ and works instead with
$\delta Q\ =\ T\,\big(\delta S\ +\ d_{i}S\big)\,$, (52)
where $d_{i}S$ is local on the patch, collects the $\mathcal{O}(\lambda^{2})$ contributions that are quadratic in the departures from equilibrium, and is non-negative in the near-equilibrium regime.
In what follows we only sketch the steps needed to extract the null-projected identity (54), referring to Ref. Eling et al. (2006) for the full non-equilibrium derivation. In a small-$\lambda$ expansion, and imposing again $\theta|_{\Sigma}=\sigma_{ab}|_{\Sigma}=0$, an appropriate choice of $d_{i}S$ (including terms of the form $\sigma^{2}$, $\theta^{2}$ and $(k\!\cdot\!\nabla f^{\prime})^{2}$) allows one to isolate the linear contributions in $\lambda$ that are sensitive to $\delta g^{ab}$. Quadratic pieces are reinterpreted as entropy production and absorbed into $d_{i}S$.
On the heat side, the calculation is identical to the area-type case: under the same assumptions on $\delta\chi^{a}$ and on the initial data, the variation of the boost charge at fixed $\beta$ still yields Eq. (44), with the same $\mathcal{O}(\lambda_{c}^{3})$ remainder.
On the entropy side, using $s(x)=\eta f^{\prime}(R)$ and expanding at small $\lambda$, one finds that the linear terms in $\lambda$ can be expressed, after an integration by parts along the generators and the use of Raychaudhuri at $\Sigma$, as a combination of
$f^{\prime}(R)\,R_{ab}k^{a}k^{b}\quad\text{and}\quad k^{a}k^{b}\nabla_{a}\nabla_{b}f^{\prime}(R)\,$, (53)
multiplied by $\eta$. Inserting these contributions into the non–equilibrium balance (52) and simplifying the common nonzero factor $\int_{-\lambda_{c}}^{0}\lambda\,d\lambda$, one obtains, using the arbitrariness of $k^{a}$ at $p$, the null-projected identity
$\frac{2\pi}{\eta}\,T_{ab}k^{a}k^{b}\ =\ \big[f^{\prime}(R)R_{ab}\ -\ \nabla_{a}\nabla_{b}f^{\prime}(R)\big]\,k^{a}k^{b}\,.$ (54)
Exactly as in the area-type branch, one completes the tensor by defining
$\mathcal{E}_{ab}\ \equiv\ f^{\prime}(R)R_{ab}\ -\ \nabla_{a}\nabla_{b}f^{\prime}(R)\ -\ \frac{2\pi}{\eta}\,T_{ab}\,$, (55)
which satisfies $\mathcal{E}_{ab}k^{a}k^{b}=0$ for every null $k^{a}$. The algebraic lemma discussed below Eq. (46) then implies $\mathcal{E}_{ab}=\Phi\,g_{ab}$ for some scalar function $\Phi$. Taking the covariant divergence, using the Bianchi identities and $\nabla^{a}T_{ab}=0$, and noting that $\nabla_{b}f^{\prime}(R)=f^{\prime\prime}(R)\nabla_{b}R$, one uniquely fixes $\Phi=\Box f^{\prime}(R)-\tfrac{1}{2}f(R)+\Lambda$, so that the full field equations read
$f^{\prime}(R)R_{ab}\ -\ \nabla_{a}\nabla_{b}f^{\prime}(R)\ +\ \big(\Box f^{\prime}(R)\ -\ \tfrac{1}{2}f(R)\big)g_{ab}\ =\ \frac{2\pi}{\eta}\,T_{ab}\,.$ (56)
Identifying $2\pi/\eta=8\pi G_{\rm eff}$, Eq. (56) coincides with the Euler–Lagrange equations obtained by varying the action
[…]
where $I_{\rm m}[g,\psi]$ is the matter action. In particular, choosing $G_{\rm eff}=G$ and $\eta=1/(4G)$ reproduces the Wald entropy density for $f(R)$ gravity Wald (1993); Iyer and Wald (1994). This shows that the canonical Massieu framework is compatible with Jacobson’s non-equilibrium extension: an area-type entropy density leads directly to Einstein’s equations with $G_{\rm eff}=1/(4s_{0})$, while a curvature-dependent density $s(x)\propto f^{\prime}(R)$ supplemented by a suitable production term $d_{i}S$ reconstructs the field equations of $f(R)$ gravity in agreement with Eling et al. (2006); Brustein and Hadad (2009); Sotiriou and Faraoni (2010); De Felice and Tsujikawa (2010). In the following, we return to the area-type branch (11) and treat $s_{0}$ as the fundamental parameter controlling the local effective coupling via Eq. (49), namely $G_{\rm eff}=1/(4s_{0})$. The remaining task is to calibrate the resolution scale $A_{*}$, and hence $s_{0}$, in a non–arbitrary way, using only intrinsic data on the horizon cross section. In the next section we formulate this requirement as a Topological Calibration Principle, based on the Gauss–Bonnet relation between area, curvature and Euler characteristic, and use it to translate generalized entropy slopes into topology- and scale-dependent constraints on admissible non-extensive exponents.
3 Topological Calibration Principle
3.1 Minimal calibration and reference area
We now examine how topology calibrates the reference scale entering the local Jacobson construction. As shown in the previous section, once the entropy slope $s_{0}$ defined in Eq. (41) is specified, horizon thermodynamics determines the effective coupling in Eq. (49) as $G_{\rm eff}=1/(4s_{0})$. From an emergent-gravity point of view, it is therefore natural to regard the area scale $A_{*}$ appearing in $s_{0}$ as fixed by the intrinsic geometry of the horizon cross section, rather than as an arbitrary external input: the local Rindler construction only uses data on the cross-section and its null congruence, and does not invoke any privileged length defined outside the patch. This motivates a minimal calibration principle for $A_{*}$, according to which the calibration area must (i) be constructed from intrinsic invariants of the cross-section, (ii) introduce no new macroscopic length scales beyond those present in the induced geometry, and (iii) remain stable under smooth deformations within a given topological class. In this section, we assume that the entropy parameters $(\eta,\delta)$ appearing in the power-law entropy (11) are universal, i.e., independent of both the horizon topology and the scale at which the slope $s_{0}$ in Eq. (41) is evaluated; whenever this assumption is relaxed (for instance by allowing $\eta(\chi)$) we will state it explicitly. For definiteness, we also restrict to the weakly non-extensive window $\tfrac{1}{2}<\delta\leq 1$ already highlighted in Sec. 2. Combining the entropy law (11) with the definition of $s_{0}$ then gives
$s_{0}\ \equiv\ s_{\delta}(A_{*})\ =\ \eta\,\delta\,(4G)^{-\delta}A_{*}^{\delta-1}\,$, (58)
so that the effective Newton constant entering the local Einstein equations becomes
$G_{\rm eff}(A_{*})\ =\ \frac{1}{4s_{\delta}(A_{*})}\ =\ \frac{G}{\eta\delta}\left(\frac{4G}{A_{*}}\right)^{\delta-1}\,.$ (59)
For $\delta=1$ and $\eta=1$ this reproduces the standard Bekenstein–Hawking values $s_{1}=1/(4G)$ and $G_{\rm eff}=G$. For $\delta\neq 1$ the same non-extensive scaling that typically modifies the standard Smarr relation shows up locally as a scale-dependent coupling, with $A_{*}$ playing the role of an RG-like “renormalization area” for the scale-dependent coupling. Formally, this is the same structural mechanism emphasized by Lü, Di Gennaro and Ong in Ref.  et al. (2025): generalized horizon entropies determine an effective Newton constant proportional to the inverse entropy slope, $G_{\rm eff}\propto s_{0}^{-1}$.
To make the geometric aspects of the calibration explicit, consider a smooth, orientable, spacelike two-surface $\Sigma$ representing a horizon cross section, with induced metric $\gamma_{ij}$, area element $\sqrt{\gamma}\,d^{2}x$, and total area
[…]
We assume that $\Sigma$ is compact and without boundary, so that the standard Gauss–Bonnet theorem can be directly applied without additional boundary terms. For such surfaces one has
$\int_{\Sigma}\!\sqrt{\gamma}\ \ {}^{(2)}\!R\ d^{2}x\ =\ 4\pi\,\chi(\Sigma)\,$, (61)
where $\chi(\Sigma)$ is the Euler characteristic of $\Sigma$ (for an orientable surface of genus $g$, $\chi(\Sigma)=2-2g$), and $^{(2)}\!R$ denotes the intrinsic scalar curvature of $(\Sigma,\gamma_{ij})$.
It is convenient to encode the intrinsic scale of a given cross-section $\Sigma$ in the area-averaged scalar curvature
$\tilde{R}_{\Sigma}\ \equiv\ \frac{1}{A}\int_{\Sigma}\!\sqrt{\gamma}\ \ {}^{(2)}\!R\ d^{2}x\,$, (62)
which equals twice the mean Gaussian curvature. In two dimensions this quantity plays a distinguished role: it is a natural global scalar with dimensions of inverse length squared, linear in $^{(2)}\!R$ and directly tied to the Gauss–Bonnet invariant. Combining Eq. (61) with the definition (62) gives
$\tilde{R}_{\Sigma}\ =\ \frac{4\pi\,\chi(\Sigma)}{A}\,$, (63)
which holds for any compact, boundaryless $\Sigma$. For $\chi(\Sigma)\neq 0$, this relation can be inverted to yield
[…]
so that, at fixed topology, specifying the area $A$ is algebraically equivalent to specifying the area-averaged intrinsic curvature $\tilde{R}_{\Sigma}$, and vice versa. We will exploit this one-to-one correspondence (for $\chi\neq 0$) to define a reference area by fixing a calibration value of $\tilde{R}_{\Sigma}$, without assuming that $^{(2)}\!R$ is constant on the physical cross section.
For spherical sections, one has $\chi=2$ and $\tilde{R}_{\Sigma}>0$, whereas for hyperbolic sections of genus $g\geq 2$, one has $\chi=2-2g<0$, and it is natural to consider negative-curvature representatives so that $\tilde{R}_{\Sigma}<0$. In both cases, the ratio $\chi/\tilde{R}_{\Sigma}$ is positive, ensuring that Eq. (64) yields a positive area. The only degenerate case is the torus, with $\chi=0$ and $\tilde{R}_{\Sigma}=0$ for flat representatives; this case will be excluded from the ratio formulas below, since the Gauss–Bonnet theorem does not fix its area at a given curvature scale.
[…]
$A_{0}(\chi)\ \equiv\ \frac{4\pi\,|\chi|}{|\tilde{R}_{*}|}\,$. (65)
[…]
$A_{*}\ \equiv\ A_{0}(\chi)\,$, (66)
so that different topologies at the same calibration scale are assigned different coarse-graining areas, and hence — via the entropy slope — different effective couplings whenever $\delta\neq 1$. One could formally allow $A_{*}=C\,A_{0}(\chi)$ with a dimensionless constant $C$, but this would reintroduce a scheme dependence in the calibration of $G_{\rm eff}$; the TCP removes this freedom by setting $C=1$.
Evaluating the entropy slope (58) at $A_{0}(\chi)$ and inserting into (59) gives the topology-dependent effective coupling
$s_{0}(\chi)\ =\ \eta\ \!\delta\ \!(4G)^{-\delta}\,A_{0}(\chi)^{\delta-1}\,,\qquad G_{\rm eff}(\chi)\ =\ \frac{1}{4s_{0}(\chi)}\ =\ \frac{G}{\eta\,\delta}\left(\frac{4G}{A_{0}(\chi)}\right)^{\delta-1}\,.$ (67)
Equation (65) can then be rearranged to yield
$G_{\rm eff}(\chi)\ =\ \frac{G}{\eta\,\delta}\left(\frac{G\,|\tilde{R}_{*}|}{\pi\,|\chi|}\right)^{\delta-1},$ (68)
[…]
$G_{\rm eff}(\chi)\ \propto\ |\chi|^{\,1-\delta}\,.$ (69)
Importantly, nothing in the local Jacobson derivation is altered: on each patch the field equations remain
$G_{ab}\ =\ 8\pi G_{\rm eff}(\chi)\,T_{ab}\ +\ \Lambda g_{ab}\,$, (70)
[…]
Although our explicit formulas have focused on the phenomenologically simple power-law branch (11), the TCP logic relies only on the existence of a well-defined entropy slope $s_{0}=\partial S/\partial A$ evaluated at the calibrated area. For a more general functional $S(A,\chi)$, potentially with explicit topological dependence, the slope $s_{0}$ computed at $A_{*}=A_{0}(\chi)$ will, in general, depend on both $A_{*}$ and $\chi$, resulting in a topology-dependent $G_{\rm eff}$ of the same structural form as in Eq. (68). In this sense, the TCP imposes a self-consistency condition on how topology may enter generalized horizon entropies: viable non-extensive laws cannot be arbitrary functions of the area alone, but must encode topological information in a manner consistent with the Gauss–Bonnet calibration.
3.2 Bounds from topology and scale
Having established Eq. (68), one can distinguish two logically independent uses, which we will now discuss separately.
3.2.1 Fixed intrinsic scale, varying topology
[…]
$\frac{G_{\rm eff}(\chi_{1})}{G_{\rm eff}(\chi_{2})}\ =\ \left(\frac{|\chi_{2}|}{|\chi_{1}|}\right)^{\delta-1}\,.$ (71)
This expresses a purely topological modulation of the locally inferred coupling at fixed intrinsic curvature scale.
As a concrete illustration, consider four–dimensional AdS black holes with horizons of constant curvature $k=\pm 1$
$ds^{2}\ =\ -f(r)\,dt^{2}\ +\ \frac{dr^{2}}{f(r)}\ +\ r^{2}\,d\Sigma_{k}^{2}\,,\qquad f(r)\ =\ k\ -\ \frac{2GM}{r}\ +\ \frac{r^{2}}{L^{2}}\,$, (72)
where $d\Sigma_{k}^{2}$ is the metric on a compact two–manifold $\Sigma_{k}$ with Euler characteristic $\chi(\Sigma_{k})$. Typical examples are $\chi=2$ for the sphere ($k=+1$) and $\chi=2-2g<0$ for higher–genus hyperbolic horizons (k=1k=-1), obtained as compact quotients of constant–curvature spaces. At the horizon $r=r_{h}$, defined as the largest positive root of $f(r)=0$, the intrinsic scalar curvature of $\Sigma_{k}$ is
$^{(2)}R\ =\ \frac{2k}{r_{h}^{2}}\,$, (73)
so that constant–curvature spherical and hyperbolic horizons with the same radius $r_{h}$ share the same intrinsic scale $|\tilde{R}_{\Sigma}|=|{}^{(2)}R|=2/r_{h}^{2}$ (for constant–curvature horizons). Evaluating the TCP at a calibration scale chosen to match this common value, $|\tilde{R}_{*}|=2/r_{h}^{2}$, Gauss–Bonnet then implies
$A_{0}(\chi)\ =\ \frac{4\pi\,|\chi|}{|\tilde{R}_{*}|}\ =\ 2\pi\,|\chi|\,r_{h}^{2}\,$, (74)
and the ratio (71) specializes to
$\frac{G_{\rm eff}(\chi_{\rm hyp})}{G_{\rm eff}(\chi_{\rm sph})}\ =\ \left(\frac{|\chi_{\rm sph}|}{|\chi_{\rm hyp}|}\right)^{\delta-1}\ =\ \left(\frac{2}{2g-2}\right)^{\delta-1}\,$, (75)
for a hyperbolic horizon of genus $g\geq 2$ compared to a spherical one.
If independent information (for instance, from bulk dynamics or holographic data) indicates that the effective coupling should not differ by more than a fractional tolerance $\Delta\in[0,1)$ between two topologies at the same intrinsic scale, one can impose
$(1+\Delta)^{-1}\ \leq\ \frac{G_{\rm eff}(\chi_{1})}{G_{\rm eff}(\chi_{2})}\ \leq\ 1+\Delta\,.$ (76)
Using (71), and the monotonicity of the logarithm, this is equivalent to the logarithmic bound
$|1-\delta|\ \leq\ \frac{\ln(1\ +\ \Delta)}{\big|\ln(|\chi_{1}|/|\chi_{2}|)\big|}\,.$ (77)
In (77) we have kept $|1-\delta|$ so that the bound is valid for both branches $\delta<1$ and $\delta>1$ and for either ordering of $|\chi_{1}|$ and $|\chi_{2}|$. For the weakly non–extensive branch $\delta\leq 1$ relevant here one simply has $|1-\delta|=1-\delta$, so the TCP constraint (77) reduces to the one-sided bound
$1\ -\ \delta\ \leq\ \frac{\ln(1\ +\ \Delta)}{\big|\ln(|\chi_{1}|/|\chi_{2}|)\big|}\,.$ (78)
This claerly shows that topology forces $\delta$ to approach the Bekenstein–Hawking value $\delta=1$ from below.
For illustration, consider a comparison between a spherical horizon ($|\chi_{\rm sph}|=2$) and a genus-$3$ hyperbolic horizon ($|\chi_{\rm hyp}|=4$). Requiring $\big|G_{\rm eff}(\chi_{\rm hyp})/G_{\rm eff}(\chi_{\rm sph})-1\big|\lesssim 10\%$ (i.e., $\Delta=0.1$) then yields
$1\ -\ \delta\ \lesssim\ \frac{\ln(1.1)}{\ln 2}\ \approx\ 0.14\,$, (79)
so $\delta\gtrsim 0.86$ in the branch $\delta\leq 1$ considered here. Taking instead a compact hyperbolic horizon of larger genus, say $g\simeq 50$ so that $|\chi_{\rm hyp}|\simeq 100$, strengthens the bound to
$1-\delta\ \lesssim\ \frac{\ln(1.1)}{\ln(50)}\ \approx\ 2\times 10^{-2}\,$, (80)
that is, $\delta\gtrsim 0.98$. These simple estimates indicate that probing a broader class of horizon topologies — such as those realized in certain AdS black hole solutions — could impose extremely stringent bounds on $1-\delta$, thereby forcing any phenomenologically viable non-extensive model to lie close to the Bekenstein–Hawking limit $\delta=1$.
It is nevertheless important to keep the status of this “purely topological” lever in proper perspective. In the observed Universe, all operationally accessible cosmological horizons appear to possess spherical topology, so that effectively only the case $\chi=2$ is directly tested; at present, the topological sector of the TCP therefore acts primarily as a conditional filter rather than as a direct observational constraint. It quantifies how strongly non-extensive exponents would be restricted if families of horizons with $|\chi|\gg 1$ were physically realized and if their effective couplings could be compared at fixed intrinsic curvature scale. Within such scenarios, Eq. (77) provides a clean bound: once $\Delta$ and the available range of $|\chi|$ are specified, sizable deviations from the Bekenstein–Hawking value become incompatible with an emergent-gravity interpretation based on horizon thermodynamics.
In holographic contexts, the TCP acquires a particularly sharp interpretation. Within AdS/CFT, the bulk Newton constant governing horizon entropy simultaneously determines the central charge (or, equivalently, the stress–tensor two-point coefficient) of the dual CFT. At fixed AdS radius, any topology dependence of $G_{\rm eff}(\chi)$ at the horizon would therefore induce a corresponding variation of the CFT central charge across distinct topological sectors. The absence of such a variation in a given holographic construction can thus be rephrased, in our framework, as a nontrivial constraint on $(\eta,\delta)$: consistency between the bulk TCP and the boundary universality of the central charge drives $\delta$ exponentially close to $1$, unless the assumption that a single non-extensive exponent governs all horizon topologies breaks down.
3.2.2 Fixed topology, running intrinsic scale
A complementary application of the TCP proceeds by keeping the topology fixed (e.g., $\chi=\chi_{0}$) and interpreting the scale dependence of $G_{\rm eff}(A_{*})$ as an RG-like running of the coarse-grained coupling as the horizon size varies within the same topological class. From Eq. (59) one then finds, assuming universal $(\eta,\delta)$, that
$\frac{G_{\rm eff}(A_{*2})}{G_{\rm eff}(A_{*1})}\ =\ \left(\frac{A_{*1}}{A_{*2}}\right)^{\delta-1}\,.$ (81)
In this second use one no longer compares different topologies: the TCP is applied within a single topological class (fixed $\chi$), while the relevant reference area is identified with the physical horizon area of the screen under consideration, $A_{*}\equiv A$ (e.g. $A\sim r_{h}^{2}$ for AdS black holes), rather than with the Gauss–Bonnet reference area $A_{0}(\chi)$ introduced for cross–topology comparisons at fixed intrinsic scale. The underlying mechanism is unchanged—the local coupling is still fixed by the entropy slope as $G_{\rm eff}(A_{*})=1/[4\,s_{\delta}(A_{*})]$—but the identification of $A_{*}$ now tracks the variation of the horizon size within the same topological sector. Any deviation from $\delta=1$ therefore appears as a controlled running of the effective coupling between, for example, small and large AdS black holes of the same topology.
In practice, the empirical near–constancy of Newton’s constant over wide ranges of curvature can be encoded by requiring that $G_{\rm eff}$ inferred from horizon thermodynamics vary by at most a fractional amount $\Delta_{\rm run}$ between two physically relevant scales. Imposing this condition on (81) gives
$(1+\Delta_{\rm run})^{-1}\ \leq\ \frac{G_{\rm eff}(A_{*2})}{G_{\rm eff}(A_{*1})}\ \leq\ 1+\Delta_{\rm run}\,$, (82)
and hence
$|1-\delta|\;\leq\;\frac{\ln(1+\Delta_{\rm run})}{\big|\ln(A_{*2}/A_{*1})\big|}\,.$ (83)
It is useful to make explicit how the constraint strengthens with the available dynamical range in area. Assuming $A_{*2}>A_{*1}$, Eq. (83) can be written as
$|1-\delta|\ \leq\ \frac{\ln(1+\Delta_{\rm run})}{\ln\!\left(A_{*2}/A_{*1}\right)}\,$, (84)
which shows that, at fixed tolerance $\Delta_{\rm run}$, the bound tightens only logarithmically as the ratio $A_{*2}/A_{*1}$ increases. For a representative choice $\Delta_{\rm run}=0.1$ and a modest dynamical range $A_{*2}/A_{*1}=100$ (for instance, two spherical AdS horizons with $r_{h,2}/r_{h,1}=10$), one finds
$|1-\delta|\ \lesssim\ \frac{\ln(1.1)}{\ln(100)}\;\approx\;2\times 10^{-2}\,$, (85)
so that $\delta\gtrsim 0.98$ in our working non–extensive window. Larger area ratios (e.g. connecting stellar–mass and supermassive horizons within the same non–extensive law) tighten the constraint according to the same logarithmic scaling.
A more extreme — and cosmologically well motivated — dynamical range within the same topological class (spherical sections, $\chi=2$) is obtained by bringing cosmological horizons into the scale-running logic of item (ii). In this extension the TCP mechanism is unchanged: $G_{\rm eff}$ is tied to the entropy slope evaluated at a resolution area $A_{*}$, which is now allowed to vary over many decades. The only additional physical input is the identification of the appropriate thermodynamic screen in a dynamical spacetime.
For an FRW universe, the standard local causal screen used in Clausius-type derivations is the apparent horizon, i.e. the marginally trapped surface on which the expansion of one of the radial null congruences vanishes (separating trapped/anti-trapped regions). It admits a geometric first–law structure closely analogous to the black–hole case Hayward (1998). Its radius is
$R_{A}(a)\ =\ \frac{1}{\sqrt{H(a)^{2}+k/a^{2}}}\ \simeq\ H(a)^{-1}\,$, (86)
Here we have assumed that $k\simeq 0$. The ensuing area is thus
$A_{H}(a)\ =\ 4\pi R_{A}^{2}\ \simeq\ \frac{4\pi}{H(a)^{2}}\,$, (87)
as in the standard apparent-horizon thermodynamic derivations of Friedmann dynamics Hayward (1998); Cai and Kim (2005). Identifying the calibration area with the apparent-horizon area, $A_{*}\equiv A_{H}$, Eq. (59) gives
$G_{\rm eff}(a)\ =\ \frac{G}{\eta\delta}\left(\frac{4G}{A_{H}(a)}\right)^{\delta-1}\ \propto\ A_{H}(a)^{\,1-\delta}\ \propto\ H(a)^{2(\delta-1)}\,$, (88)
where the omitted prefactor is $(G/(\eta\delta))\,(G/\pi)^{\delta-1}$ in the $k\simeq 0$ approximation, using $A_{H}(a)\simeq 4\pi/H(a)^{2}$. Since our focus is the running, it is convenient to remove the scheme-dependent prefactor by working with the dimensionless modulation normalized to unity today
$\mu(a)\ \equiv\ \frac{G_{\rm eff}(a)}{G_{\rm eff}(1)}\ =\ \left[\frac{A_{H}(a)}{A_{H}(1)}\right]^{1-\delta}\ =\ \left[\frac{H(a)}{H_{0}}\right]^{2(\delta-1)}\,.$ (89)
Thus, under the universality assumption for $(\eta,\delta)$ and the identification $A_{*}=A_{H}$, the TCP yields a definite redshift scaling for the effective gravitational coupling (given an assumed background expansion history $H(a)$): the time dependence of $\mu(a)$ is fixed by $H(a)$, with the exponent given directly by the non-extensivity index $\delta$.
A first diagnostic follows by applying the scale-running bound (83) to a deliberately conservative comparison between the present apparent-horizon area and a supermassive black-hole horizon area. Taking $R_{A0}\equiv R_{A}(a{=}1)\sim H_{0}^{-1}$ and $R_{s}\sim 2GM$ (up to $\mathcal{O}(1)$ spin-dependent factors, irrelevant at the level of this estimate) gives $A_{H}/A_{\rm BH}\sim(R_{A0}/R_{s})^{2}\sim 10^{27\text{--}28}$ at the order-of-magnitude level, hence $\log_{10}(A_{H}/A_{\rm BH})\sim 27$2828. Requiring a fractional stability $|\Delta G/G|\lesssim\Delta_{\rm run}$ across this range yields
$|1-\delta|\ \lesssim\ \frac{\ln(1+\Delta_{\rm run})}{\big|\ln(A_{H}/A_{\rm BH})\big|}\ \sim\ 10^{-3}\,,\qquad\text{for}\ \ \Delta_{\rm run}\sim 10^{-1}\,.$ (90)
This estimate should be read as a consistency filter under the working identification of the thermodynamic coupling with the effective gravitational strength constrained by astrophysical and cosmological probes. For percent-level tolerances, $\Delta_{\rm run}\sim 10^{-2}$—as suggested by existing bounds on $|\Delta G/G|$ from early- and late-universe consistency tests—the constraint tightens to $|1-\delta|\lesssim 2\times 10^{-4}$ Uzan (2011). In particular, benchmark values such as the microcanonical exponent $\delta=3/2$ would imply $G_{\rm eff}(A_{H})/G_{\rm eff}(A_{\rm BH})\sim(A_{H}/A_{\rm BH})^{1/2}\sim 10^{13\text{--}14}$, incompatible by many orders of magnitude unless the universality assumption on $\delta$ is abandoned.
[…]
$D^{\prime\prime}(a)\ +\ \left[\frac{3}{a}\ +\ \frac{H^{\prime}(a)}{H(a)}\right]D^{\prime}(a)\ -\ \frac{3}{2}\frac{\Omega_{m}(a)}{a^{2}}\,\mu(a)\,D(a)\ =\ 0\,$, (91)
[…]
A dedicated confrontation of the specific redshift scaling (89) with survey likelihoods would require embedding the TCP prediction into the growth-sector parameterizations used in full-shape and weak-lensing pipelines, including consistent treatment of bias and projection effects. Forecast validations for upcoming surveys indicate percent-level sensitivity to scale-independent growth modifications over $0\lesssim z\lesssim 2$, making (89) a concrete target for next-generation large-scale-structure data Collaboration et al. (2020); Collaboration and others (2022). In summary, applying the TCP scale-running logic to cosmology yields two concrete outputs: a stringent consistency bound (90) that forces any universal non–extensive exponent extremely close to the Bekenstein–Hawking value $\delta=1$, and a falsifiable growth signature (89)–(91) whose characteristic redshift dependence can be targeted by ongoing and forthcoming surveys.
3.3 Status and variants of the TCP
[…]
$G_{\rm eff}(\chi)\ \equiv\ G_{\rm phys}\,$, (92)
for all admissible surfaces (spheres, higher-genus hyperbolic sections, and so on). With universal $(\eta,\delta)$ this immediately forces $\delta=1$, so that the non–extensive branch collapses back to the Bekenstein–Hawking law. If one wants to keep $\delta\neq 1$ while enforcing $G_{\rm eff}=G_{\rm phys}$, one must then relax the universality of $\eta$ and allow it to track topology. Solving (68) for $\eta$ at fixed $\delta$ and $G_{\rm phys}$ gives
$\eta(\chi)\ =\ \frac{G}{\delta\,G_{\rm phys}}\left(\frac{G\,|\tilde{R}_{*}|}{\pi\,|\chi|}\right)^{\delta-1}.$ (93)
In this “non–extensive but universal” variant the Jacobson equations on each patch are still Einstein with coupling $G_{\rm phys}$, but the global entropy prefactor ceases to be universal and instead carries topological information. A simple example of this logic is provided by Einstein gravity supplemented by a four-dimensional Gauss–Bonnet term: in that case the Wald entropy acquires an additive contribution $S_{\rm GB}\propto\alpha\,\chi(\Sigma)$ with vanishing slope $dS_{\rm GB}/dA=0$, so that the Gauss–Bonnet piece acts as a purely topological counterterm which can be used to calibrate the overall normalization of $S(A)$ without modifying the local field equations; the corresponding algebraic relations between $(\eta,\delta)$ and the coupling $\alpha$ are collected in Appendix A.
[…]
It is therefore useful to summarize explicitly the hypothesis space within which our quantitative bounds should be interpreted.
  1. 1.
    We assume that horizon entropies are of area type, $S(A)=\eta(A/4G)^{\delta}$, with a well-defined slope $s_{0}=\partial S/\partial A$ that determines a thermodynamic coupling $G_{\rm eff}=1/(4s_{0})$ as in Eq. (59).
  2. 2.
    When comparing different topologies at fixed intrinsic curvature scale, we impose the TCP calibration $A_{*}=A_{0}(\chi)$, thereby implementing the “no external scales” requirement discussed in Sec. 3.1; for scale-running arguments at fixed topology, we instead identify $A_{*}$ with the physical horizon area.
  3. 3.
    Whenever $G_{\rm eff}$ is confronted with data, we adopt the working hypothesis that the thermodynamic coupling inferred from local horizon patches can be identified, up to order-unity scheme ambiguities, with the effective Newton constant $G_{\rm phys}$ constrained by laboratory, astrophysical, and cosmological measurements.
Within this framework, the TCP introduces no additional dynamics beyond the local Jacobson construction; rather, it acts as a calibration principle that renders non-extensive area-type proposals quantitatively testable. Once $(\eta,\delta)$ are specified, topology and scale feed unavoidably into $G_{\rm eff}$ through Gauss–Bonnet theorem and the Clausius relation, so that resulting variation of the effective coupling across horizons becomes a sharp target for both microscopic models and phenomenological bounds.
4 Conclusion and outlook
In this work we have combined non-extensive horizon thermodynamics with Jacobson’s emergent-gravity paradigm to investigate how generalized area-type entropies can be incorporated while maintaining a transparent thermodynamic origin of the field equations. Our results exhibit two complementary layers: a local layer, where the dynamics is derived on Rindler patches in canonical form, and a global layer, where intrinsic geometry and topology calibrate the evaluation scale entering the local construction.
At the local level, we recast Jacobson’s horizon argument as the stationarity of a Massieu functional at fixed Unruh temperature. In the area-type sector—where the entropy density is uniform on the patch and determined by the slope $s_{0}=dS/dA|_{A_{*}}$ of a phenomenological power law $S(A)=\eta(A/4G)^{\delta}$ — the canonical balance reproduces Einstein’s field equations with an effective Newton coupling $G_{\rm eff}=1/(4s_{0})$. Allowing instead for curvature dependence, $s(x)\propto f^{\prime}(R)$, and including an internal entropy-production term in the sense of Jacobson’s non-equilibrium extension, the same framework yields the field equations of $f(R)$ gravity. In both regimes, the conclusion is clear: the only macroscopic input probed locally by the horizon balance is the entropy slope evaluated at the chosen reference area $A_{*}$.
[…]
These two applications yield explicit logarithmic constraints on $|1-\delta|$. At fixed intrinsic scale, comparing two Euler characteristics leads to the bound (77); at fixed topology, comparing two areas yields the running bound (83). When $G_{\rm eff}$ is identified with the effective gravitational coupling constrained by observational data, the enormous dynamical range between black-hole and cosmological horizon areas turns the scale-running bound into a stringent consistency filter, forcing any universal non-extensive exponent to lie extremely close to the Bekenstein–Hawking value $\delta=1$. In particular, benchmark values suggested by microcanonical group-entropy arguments (such as $\delta=3/2$) would imply variations of $G_{\rm eff}$ across horizons that are far too large, unless universality across sectors is relaxed or the identification between the local thermodynamic coupling and the Newtonian limit is modified.
A notable byproduct of the TCP is that it provides concrete, falsifiable signatures rather than merely consistency bounds. In cosmology, identifying the apparent-horizon area as the evaluation scale for the entropy slope, $A_{*}\equiv A_{H}\simeq 4\pi/H^{2}$, leads to a fixed redshift dependence for the normalized coupling modulation, $\mu(a)=G_{\rm eff}(a)/G_{\rm eff}(1)=[H(a)/H_{0}]^{2(\delta-1)}$, cf. Eq. (89). Propagating this prediction into the standard linear growth equation (while keeping the background expansion fixed) produces a characteristic imprint on $f\sigma_{8}(z)$, depicted in Fig. 2. The essential point is structural: within the TCP framework, the growth-sector modification is not an arbitrary function but a one-parameter deformation whose redshift dependence is fixed by the observed expansion history. Large-scale-structure measurements therefore provide a natural arena in which to test the TCP-induced running in a minimally model-dependent manner.
From a microscopic perspective, the TCP can be viewed as a macroscopic selection rule. Any candidate quantum-gravity mechanism yielding a generalized horizon entropy must not only reproduce an admissible local slope $dS/dA$ consistent with Jacobson’s balance, but also account for the near invariance of the corresponding thermodynamic coupling across the enormous range of horizon scales and, where relevant, across topological sectors. In this sense, topology is not merely a descriptive label: through the Gauss–Bonnet relation it becomes a quantitative probe of how generalized entropies may encode intrinsic geometric information without reintroducing external scales.
Several natural extensions our work suggest themselves. First, it would be worthwhile to incorporate the TCP running $\mu(a)$ into full cosmological inference procedure — including weak-lensing and full-shape analyses — to quantify the sensitivity to $|1-\delta|$. Second, holographic frameworks offer a particularly intriguing theoretical arena: because the bulk Newton constant fixes boundary central charges, any TCP-induced modulation of $G_{\rm eff}$ across horizon topologies should leave a boundary imprint, potentially elevating the topology lever to a precision consistency test. Third, extending the calibration logic beyond $f(R)$ to more general higher-curvature gravity theories (e.g. Lanczos–Lovelock or Weyl) and to entropy functionals with explicit dependence on curvature invariants may clarify the extent to which the “slope controls coupling” mechanism persists when the horizon entropy becomes genuinely non-local or curvature sensitive.
Overall, the picture that emerges is that horizon thermodynamics not only accommodates non-extensive proposals but also constrains them. Once the local Clausius balance is taken seriously, the entropy slope becomes the operationally relevant quantity, while topology provides a principled means of calibrating the associated area scale. Taken together, these ingredients render generalized area-type entropies quantitatively constrained and observationally testable deformations of the Bekenstein–Hawking law.
Acknowledgements – G.L. acknowledges the Istituto Nazionale di Alta Matematica (INdAM), Gruppo Nazionale di Fisica Matematica. P.J. acknowledges support from the Czech Science Foundation Grant Agency (GAČR), Grant No. 25-18105S.
Appendix A Gauss–Bonnet entropy as a topological counterterm
In this appendix we illustrate how the calibration logic discussed in Sec. 3 operates in a theory where the gravitational Lagrangian already contains a purely topological density. A standard example is Einstein gravity supplemented by the four-dimensional Gauss–Bonnet density
$\mathcal{L}\ =\ \frac{1}{16\pi G}\,\big(R\ +\ \alpha\,\mathcal{G}\big)\,,\qquad\mathcal{G}\ =\ R_{abcd}R^{abcd}\ -\ 4R_{ab}R^{ab}\ +\ R^{2}\,$, (94)
where $\mathcal{G}$ is the quadratic Gauss–Bonnet term. In four dimensions its integral is proportional to the Euler characteristic, so the equations of motion remain those of Einstein gravity while the Wald entropy acquires an additive topological contribution Wald (1993); Iyer and Wald (1994)
$S_{\rm GB}\ =\ \frac{2\pi\,\alpha}{G}\,\chi(\Sigma)\,.$ (95)
In particular, $S_{\rm GB}$ provides exactly the kind of slope-free topological counterterm envisaged by the TCP: since $dS_{\rm GB}/dA=0$, it shifts the entropy zero point and its global topological dependence without altering the slope $s_{0}$ and hence the local Jacobson equations.
In this setting the TCP logic can be used in a slightly different way: the Gauss–Bonnet contribution calibrates the total entropy at a fiducial area without modifying the local field equations. Concretely, define
$S_{\rm tot}(A,\chi)\ =\ S(A)\ +\ S_{\rm GB}(\chi)\ =\ \eta\Big(\frac{A}{4G}\Big)^{\delta}\ +\ \frac{2\pi\alpha}{G}\,\chi\,$, (96)
and choose a reference spherical section with $\chi=2$ and area $A_{\rm ren}$, naturally tied to the intrinsic curvature scale by $A_{\rm ren}=A_{0}(2)=8\pi/\tilde{R}_{*}$, where $A_{0}(\chi)$ is the Gauss–Bonnet reference area introduced in Eq. (65). Imposing slope matching to the physical Newton coupling,
$\frac{dS_{\rm tot}}{dA}\Big|_{A_{\rm ren}}\ =\ \frac{1}{4\,G_{\rm phys}}\,$, (97)
and area-law normalization on the sphere
$S_{\rm tot}(A_{\rm ren};\chi=2)\ =\ \frac{A_{\rm ren}}{4\,G_{\rm phys}}\,$, (98)
one finds
$\eta\ =\ \frac{G}{\delta\,G_{\rm phys}}\left(\frac{A_{\rm ren}}{4G}\right)^{1-\delta}\,,\qquad\alpha\ =\ \frac{G}{16\pi}\,\frac{A_{\rm ren}}{G_{\rm phys}}\left(1\ -\ \frac{1}{\delta}\right)\,.$ (99)
Equivalently, for given $(\alpha,G_{\rm phys})$ the calibration fixes $\delta$ and hence $\eta$. This example shows explicitly how the TCP logic can also relate non-extensive entropy parameters to a topological Gauss–Bonnet coupling in theories whose local dynamics remain Einsteinian.